OpenStudy (anonymous):

State the limits of the following sequences, or state that the limit does not exist 5, 6 1/2, 5 2/3, 6 1/4, 5 4/5, 6 1/6,..., 6+(-1)^n/n.. Can anyone explain why that the answer is \[\lim_{n \rightarrow \infty}6+(-1)^n/n = 6\]?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

the larger \(n\) gets the close \(\frac{(-1)^n}{n}\) gets to zero

OpenStudy (anonymous):

So it means \[\lim_{n \rightarrow \infty} 6+(0) = 6?\]

OpenStudy (misty1212):

\[\lim_{n \rightarrow \infty} 6+\frac{(-1)^n}{n}\] \[=6+\lim_{n\to \infty}\frac{(-1)^n}{n}=6+0=6\]

OpenStudy (anonymous):

okay thanks!

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]